3x - 2y = 11
x - y/2 = 4
how do i solve these equations using substation?

Answer by checkley77(12844)

3x - 2y = 11
x - y/2 = 4 multiply by 3 & subtract.
3x-2y=11
-(3x-3y/2=12)
----------------
-2y+3y/2=-1
(-2y*2+3y)/2=-1
(-4y+3y)/2=-1
-y/2=-1
-y=-2
y=2 ans.
3x-2*2=11
3x-4=11
3x=11+4
3x=15
x=15/3
x=5
Proof:
5-2/2=4
5-1=4
4=4